OpenStudy (anonymous):

Please help me integrating this:

OpenStudy (anonymous):

\[\int\limits_{?}^{?} cosx dx / sinx (2+sinx)\]

OpenStudy (anonymous):

it's an indefinite integral

OpenStudy (anonymous):

help me!!!!!!!

OpenStudy (watchmath):

Use substitution \(u=\sin x\)

OpenStudy (anonymous):

hey watch, do you know how to view old questions, in your settings ?

OpenStudy (watchmath):

old questions that you asked or old questions that you answered?

OpenStudy (anonymous):

for instance that induction problem we were working on, is lost

OpenStudy (watchmath):

I can see it, by the link is related to my profile. Try to click the link here : http://openstudy.com/users/watchmath#/users/watchmath/updates/4dd9bc1c0ada8b0b731c42c7

OpenStudy (anonymous):

so did that question make sense earlier

OpenStudy (anonymous):

the product of e^(r/n) *1/n = y , they took the log of both sides

OpenStudy (anonymous):

it isnt quite exactly the integral since i thought you need a right hand or left hand sum

OpenStudy (anonymous):

ln y = 1/n [ e^0/n + e^(1/n) + ... e^n/n)

OpenStudy (anonymous):

the right hand side is not quite a riemann sum,

OpenStudy (anonymous):

lim n-> oo 1/n [ e^0/n + e^(1/n) + ... e^n/n)

OpenStudy (watchmath):

Let's continue on the old thread: http://openstudy.com/users/watchmath#/users/watchmath/updates/4dda53c30ada8b0b11c243c7